OpenStudy (anonymous):

Choose the correct product of (5x − 3)(5x + 3)

OpenStudy (hba):

It's very easy to multiply this..

OpenStudy (aravindg):

(a+b)(a-b) =?

OpenStudy (theviper):

\[a^2-b^2\]

OpenStudy (aravindg):

exactly use that and u are done :)

OpenStudy (hba):

You can use the formula if you don't wanna use FOIL

OpenStudy (hba):

(a+b)(a-b)=(a)^2-(b)^2

OpenStudy (anonymous):

okay so i foil it ?

OpenStudy (hba):

So if here a=5x b=3 then

OpenStudy (hba):

@andrewsurfs4995 You can also foil it but using the formula is an easy method.

OpenStudy (theviper):

Yes @andrewsurfs4995 \[(5x − 3)(5x + 3)=(5x)^2-(3)^2\]